An Oil Tanker can be filled in 10 hrs from one pipe. It can be filled by another pipeline in 8 hrs. How long will it take to fill the tank if both pipe lines are used.
answer: it will take 4 4/9 hoursto fill the tanker using both pipes.

need a break down of how my book got 4 4/9 hours to fill the tanker using both pipes.

Answer by checkley77(12844)

10*8/(10+8)=80/18=4.444 HOURS TO FILL THE TANK USING BOTH PIPES.